OpenStudy (anonymous):

Use a half angle identity to find the approximate value of sin15(degrees)

OpenStudy (anonymous):

half of 30 is 15 use the half angle formula with \(\theta=30\) and \(\frac{\theta}{2}=15\)

OpenStudy (anonymous):

A. 0.26 b. 0.5 c. 1 d. radical 2

OpenStudy (anonymous):

god bless multiple choice

OpenStudy (anonymous):

since \(\sqrt2>1\) you can ignore that one

OpenStudy (anonymous):

since \(\sin(90)=1\) you can ignore that one too

OpenStudy (anonymous):

either a or b?

OpenStudy (anonymous):

do you know the half angle identity?

OpenStudy (anonymous):